# iRAG: An Incremental Retrieval Augmented Generation System for Videos

### v1

@@ V Prototype Implementation
In this example, the DETR object detection model is employed for preprocessing. Now, for detection-based queries such as “Is there a bus?”, the indexer efficiently indexes the video and provides the query’s answer. For more advanced user queries, like “What is the color of the bus?”, the indexer suggests the necessary video clips. The Extractor then uses the GRiT model to extract detailed text information from these clips. Subsequently, the extracted text serves as context, and the LLM API is invoked along with the query to generate the final response.
VI Related Work
Traditional Video Understanding Models. Traditional video understanding models have focused on designing foundational models for specific tasks [20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. These models distinguished themselves by incorporating a variety of neural network architectures and employing diverse training methodologies. These models demonstrate superior performance when analyzing short video clips lasting several seconds, but they encounter difficulty when tasked with processing longer videos.
Video-Language Models. Researchers have made significant progress in the field of multimodal representation learning [30, 31, 32], particularly emphasizing the convergence of video and language understanding [33, 34, 35]. They employ a two-stage approach, commencing with the pretraining of a foundational video-language model using extensive pairs of video and text, and subsequently fine-tuning on specific downstream tasks. Similarly, recent studies propose the integration of a frozen Language-Modeling Module (LLM) alongside an image encoder and a limited set of learnable modules, introducing a technique termed “visual instruction tuning” [36, 37]. The main goal of this technique is to fine-tune the learnable modules, aiming to enhance the LLMs’ capacity to generate textual descriptions for video content. Although all these models exhibit promising performance, they may face challenges when dealing with long-duration videos. Notably, these models are typically trained on short video clips, potentially introducing difficulties in understanding longer video sequences.
@@ VI Related Work
Video question answering system [38, 39, 40] is extensively studied in the literature. Unfortunately, most of them mainly focus on short video clips of images. In recent works, the focus has been on understanding complex relationships in long videos, such as cause-and-effect and sequential order [41, 42]. Typically, questions in long-form video question-answering systems require reasoning about multiple events, various levels of detail, and causality. For each query, with efficient segmentation of videos, the whole work might be repetitive for searching the corresponding relevant frames which is almost impossible for long videos.
LLMs in Text-Based Video Understanding. Large language models (LLMs) like ChatGPT have exhibited notable proficiency in handling text-based dialogues, and effectively managing input and output in textual format. Recently, LLMs connect visual foundation models that create a new dimension for advanced video understanding [43, 44]. Inspired by the success of reasoning and interaction with LLMs, recent studies have redefined the challenge of understanding long video content through a question-and-answer framework [4, 5, 35]. Initially, these methods segment a lengthy video into smaller clips and utilize readily available vision and vision-language models to convert visual data from each clip into text. Subsequently, they employ LLMs to generate textual responses to queries associated with the video, utilizing textual information as context. In contrast to these methods, which typically rely on a set of models to extract textual information from videos upfront, our approach aims to execute models incrementally based on queries, extracting the necessary textual information on-demand to address each specific query.
VII Conclusion
We introduced a novel incremental Retrieval-Augmented Generation (iRAG) system for enhancing understanding of multimodal data like videos and images.
Instead of using many AI models to upfront extract text data from the entire corpus of multimodal data, iRAG first quickly indexes the corpus of multimodal data by using lightweight AI models. There will be user queries for which the information extracted from the multimodal data during index building is insufficient. In such cases,
@@ VII Conclusion
iRAG invokes the incremental workflow to extract more detailed information from select, query relevant multimodal data. Such new information is used to retrieve an updated context, which results in a useful response to the user query. We evaluated iRAG for interactive video understanding, but the iRAG system is useful for combined understanding of a variety of other multimodal data like audio, lidar and time-series. The AI models required to capture content in such multimodal data as text, or quickly build indexes, are different. However, like video, such multimodal data also has time as a marker, and the multimodal content is easily embedded into vectors, making iRAG indispensable for advanced, interactive understanding of multimodal data.

### v2

@@ 4. Prototype Implementation
In this example, the DETR object detection model (Carion et al., 2020) is employed for preprocessing. Now, for detection-based queries such as “Is there a bus?”, the indexer efficiently indexes the video and provides the query’s answer. For more advanced user queries, like “What is the color of the bus?”, the indexer suggests the necessary video clips. The Extractor then uses the GRiT model (Nguyen et al., 2022) to extract detailed text information from these clips. Subsequently, the extracted text serves as context, and the LLM API is invoked along with the query to generate the final response.
Figure 7. iRAG prototype implementation
5. Related Work
Traditional video understanding models have primarily concentrated on designing task-specific models, incorporating various neural network architectures and training techniques, and excelling in the analysis of short video clips, as evidenced by previous works (Lin et al., 2019; Feichtenhofer et al., 2019; Bertasius et al., 2021; Zhang et al., 2021; Arnab et al., 2021; Li et al., 2022; Debnath et al., 2022). However, this paper shifts its focus to analyzing long videos.
In recent years, significant progress has been made in multimodal learning, particularly in integrating video and language understanding, as demonstrated by recent research (Li et al., 2023a, b; Maaz et al., 2023). Researchers have explored techniques like “visual instruction tuning”, which combines a fixed Language Model (LLM) with a vision model and adaptable modules to refine LLMs’ ability to generate textual descriptions for video content (Liu et al., 2023; Zhang et al., 2023b). While these models are utilized in systems for answering questions about videos (Wang et al., 2023; Hu et al., 2023; Zhang et al., 2023a; Arefeen et al., 2024b), most of them are tailored for short video clips or images. This paper, however, aims to improve the runtime efficiency of the video-to-text extraction process specifically for long videos within a retrieval-augmented generation (RAG) (Lewis et al., 2020) based system.
@@ 5. Related Work
Video question answering systems (Wang et al., 2023; Hu et al., 2023; Zhang et al., 2023a) have been extensively researched, but primarily on short video clips or images. Recent works, however, have shifted focus to understanding complex relationships in long videos, such as cause-and-effect and sequential order (Gao et al., 2021, 2023). Long-form video question-answering systems typically involve reasoning about multiple events, various levels of detail, and causality. Efficiently segmenting videos for each query can be repetitive and challenging, especially for long videos.
6. Conclusion
@@
We introduced an innovative incremental Retrieval-augmented Generation (iRAG) system to enhance the understanding of videos. Instead of using many AI models to extract text data from the entire video corpus upfront, iRAG quickly indexes the video data using lightweight AI models. When the extracted information during indexing is insufficient for certain user queries, iRAG employs an incremental workflow to extract more detailed information from select, query-relevant video data using heavyweight AI models. This new information is then used to retrieve an updated context, resulting in a useful response from the LLM to user query. While we evaluated iRAG for interactive video understanding, it is applicable to a variety of other non-textual data types like audio, LiDAR, and time-series data. The required AI models for such data differ, but iRAG remains indispensable due to its ability to handle time-sensitive non-textual content efficiently.